# Estimating non-linear functionals of trawl processes

### v2

@@ 4.3 Proof of Theorems 5 and 6
Next, let $\tilde{m}_{n}=[T/\Delta_{n}]$ if $c_{l}^{n}(t)=\mathbf{1}_{0\leq l\leq[t/\Delta_{n}]-1}$ and $\tilde{m}_{n}=m_{n}$ otherwise and note that Assumption 1 implies that the estimate in (42) is valid. Thus,
$\sup_{t\in[0,T]}\rvert U_{t}^{n,2}\rvert\lesssim\sqrt{n\Delta_{n}^{3}}\int_{0}^{\tilde{m}_{n}\Delta_{n}}\dot{g}(a(\tau_{n}(s)))\phi(s)\mathrm{d}s.$
Since $\sqrt{n\Delta_{n}^{3}}\rightarrow 0,$ and $\phi$ is bounded ($\phi$ is càdlàg and vanishes at $\infty$), to see that the result holds for $i=2,$ we only need to verify that under our assumptions the integral $\int_{0}^{m_{n}\Delta_{n}}\dot{g}[a(\tau_{n}(s))]\mathrm{d}s$ is bounded over $n$. Indeed, if $\tilde{m}_{n}=[T/\Delta_{n}]$ the statement follows trivially by the continuity of $\dot{g}$. If $\tilde{m}_{n}=m_{n}$ and $g\in\mathfrak{C}_{p,q}^{2}$, Remark 1 implies that
$\int_{0}^{m_{n}\Delta_{n}}\rvert\dot{g}[a(\tau_{n}(s))]\rvert\mathrm{d}s\lesssim\int_{0}^{m_{n}\Delta_{n}}a(\tau_{n}(s))^{1+p}\mathrm{d}s\lesssim\Delta_{n}+\int_{0}^{\infty}a(s)^{1+p}\mathrm{d}s,$ (54)
just as required. For $i=3$, we observe that (recall that $xa(x)\rightarrow 0$)
$\sup_{t\in[0,T]}\rvert U_{t}^{n,3}\rvert\lesssim_{a}\frac{1}{\sqrt{n\Delta_{n}}}\int_{0}^{\tilde{m}_{n}\Delta_{n}}\dot{g}[a(\tau_{n}(s))]\mathrm{d}s\rightarrow 0,$
due to (54). Similar arguments let us infer that $\sup_{t\in[0,T]}\rvert U_{t}^{n,4}\rvert\overset{\mathbb{P}}{\rightarrow}0$. To deal with the final case suppose first that $c_{l}^{n}(t)=\mathbf{1}_{0\leq l\leq[t/\Delta_{n}]-1}$. Thus, in view that $g$ of class $\mathcal{C}^{2}$ and $a$ is Lipschitz (thanks to Assumption 1), we deduce that for all $t\in[0,T]$
$\sqrt{n\Delta_{n}}\rvert\mathcal{R}_{t}^{n}\rvert$ $\lesssim\sqrt{n\Delta_{n}}\int_{0}^{[t/\Delta_{n}]\Delta_{n}}\left|g(a(\tau_{n}(s)))-g(a(s))\right|\mathrm{d}s+\sqrt{n\Delta_{n}^{3}}$
$\lesssim\sqrt{n\Delta_{n}^{3}}\rightarrow 0.$
Similarly, if either $c_{l}^{n}(t)=\mathbf{1}_{[t/\Delta_{n}]\leq l\leq n-1}$ or $c_{l}^{n}(t)=\mathbf{1}_{[t/\Delta_{n}]\leq l\leq N_{n}-1}$ and $g\in\mathfrak{C}_{p,q}^{2},$ we get from Remark 1 that
$\rvert\mathcal{R}_{t}^{n}\rvert\lesssim\int_{0}^{m_{n}\Delta_{n}}\rvert g(a(\tau_{n}(s)))-g(a(s))\rvert\mathrm{d}s+\int_{m_{n}\Delta_{n}}^{\infty}\rvert g(a(s))\rvert\mathrm{d}s+\Delta_{n}.$
Moreover, from (11) we deduce that $\rvert g(a(s))\rvert\leq Ca(s)^{2+p}$ for all $s>0$ and by the Mean-Value Theorem and the Lipshchitz condition, there is $x_{s,n}\in[a(s),a(\tau_{n}(s))]$, such that
$\rvert g(a(\tau_{n}(s)))-g(a(s))\rvert\lesssim\rvert\dot{g}(x_{s,n})\rvert\Delta_{n}\lesssim\Delta_{n}a(\tau_{n}(s))^{p+1},$
where in the second inequality we used also (11). A further application of Assumption 1 and (54) along with the previous estimates allow us to conclude that for $n$ large, uniformly on $t\in[0,T]$
@@
Therefore, if $m_{n}=n$, we have $\sqrt{n\Delta_{n}}(m_{n}\Delta_{n})^{1-\alpha(p+2)}=(n\Delta_{n})^{\frac{3}{2}-\alpha(p+2)}\rightarrow 0$ because $\alpha>1$. In contrast, if $m_{n}=N_{n}$ and Assumption 2 holds with $\kappa$ as in (16), we have that
@@
The next step is to write $V^{n}$ defined in (52) as a sum of of martingale differences plus an error term. Once we are done with this, the proof of Theorems 5 and 6 will follow as an application of e.g. Theorem 6.1 in [8] (c.f. Theorem IX 7.19 in [12]). We do this by means of the decomposition (32):
@@
where
$V_{t}^{\prime n}$= $\sqrt{\frac{\Delta_{n}}{n}}\sum_{l=0}^{n-1}\sum_{j=l}^{n-1}c_{l}^{n}(t)\dot{g}(a(l\Delta_{n}))\varXi_{j,l,n};$ (55)
$U_{t}^{\prime,n}$= $\sqrt{\frac{\Delta_{n}}{n}}\sum_{l=0}^{n-1}\sum_{j=l}^{n-1}c_{l}^{n}(t)\dot{g}(a(l\Delta_{n}))R_{j,l,n}.$
Clearly $\varXi_{j,l,n}$ is $\mathscr{G}_{j}^{n}$-measurable (see (23)) and by construction $\mathbb{E}(\varXi_{j,l,n}\rvert\mathscr{G}_{j-1}^{n})=0$, so is a $(\mathscr{G}_{j}^{n})_{j\geq 0}$ martingale difference. Let us now see that $U^{\prime,n}$ is an error term.
@@
As before we may and do assume that $L$ has mean 0. Since each $\rho_{j,l,n}^{(\ell)}$ (see (34)) is a martingale difference, we only need to show that for $\ell=1,2,3$
$\frac{1}{n\Delta_{n}}\mathcal{E}_{n}^{(\ell)}(t):=\frac{1}{n\Delta_{n}}\sum_{j=0}^{n}\mathbb{E}\left[\left(\Delta_{n}\sum_{l=0}^{j}c_{l}^{n}(t)\dot{g}(a(l\Delta_{n}))\rho_{j,l,n}^{(\ell)}\right)^{2}\right]\rightarrow 0.$
To see that this is certainly the case, first note that $\Delta_{n}\sum_{l=0}^{j}\rvert c_{l}^{n}(t)\dot{g}(a(l\Delta_{n}))\rvert$ is uniformly bounded. Indeed, if $c_{l}^{n}(t)=\mathbf{1}_{0\leq l\leq[t/\Delta_{n}]-1}$, by the continuity of $\dot{g}$ and $a$
$\Delta_{n}\sum_{l=0}^{j}\rvert c_{l}^{n}(t)\dot{g}(a(l\Delta_{n}))\rvert\leq\Delta_{n}\sum_{l=0}^{[T/\Delta_{n}]}\lvert\dot{g}(a(l\Delta_{n}))\rvert\leq CT.$
If instead either $c_{l}^{n}(t)=\mathbf{1}_{[t/\Delta_{n}]\leq l\leq n-1}$ or $c_{l}^{n}(t)=\mathbf{1}_{[t/\Delta_{n}]\leq l\leq N_{n}-1}$ and $g\in\mathfrak{C}_{p,q}^{2},$ we deduce from (54) that our statement is also true in this case. Hence, by Jensen’s inequality
$\mathcal{E}_{n}^{(\ell)}(t)\lesssim\Delta_{n}\sum_{j=0}^{n}\sum_{l=0}^{j}\rvert c_{l}^{n}(t)\dot{g}(a(l\Delta_{n}))\rvert\mathbb{E}[(\rho_{j,l,n}^{(\ell)})^{2}].$
Furthermore, by Lemma 1
@@
and by the Dominated Convergence Theorem we also have that
@@
We are now ready to present a proof for the CLT associated to our statistics. Note that in view of the previous two lemmas and the definition of $\tilde{a}$ (see (26)), we may and do assume that $L$ has mean 0 for the rest of this work.
@@
We start making some observations. As pointed out above, we only need to concentrate on $V^{n}$ (see (52)). Note also that thanks to Lemma 3, for all $0\leq s\leq t\leq T$ we have
@@
Therefore, according to [5] (Theorem 13.5 and equation (13.14)), Remark 2, and Lemma 6, we only need to show that the finite-dimensional distributions of $V^{\prime n}$ (for its definition see (55)) converge to those of $Z$, with $Z$ as in the theorems. We will use the Cramér-Wold methodology. Fix $M\in\mathbb{N}$ and let $0=t_{0}\leq t_{1}\leq\cdots\leq t_{M}\leq T$ and $z_{1},z_{2},\ldots,z_{M}\in\mathbb{R}$. By setting $\overline{c}_{l}^{n}=\sum_{k=1}^{M}z_{k}c_{l}^{n}(t_{k})\dot{g}(a(l\Delta_{n})),$ and
$\zeta_{j}^{n}=\Delta_{n}\sum_{l=0}^{j}\overline{c}_{l}^{n}\varXi_{j,l,n},$ (56)
in which $\varXi_{j,l,n}=\sum_{\ell=1}^{4}\xi_{j,l,n}^{(\ell)}$ (see (33)) we have that
@@
As pointed out above, $\zeta_{j}^{n}$ is a $(\mathscr{G}_{j}^{n})_{j\geq 0}$-martingale difference. Therefore, according to Theorem 6.1 in [8] (c.f. Theorem IX 7.19 in [12]), we need to show that for some $r>2$
$1.\,\sigma_{n}^{2}:=\frac{1}{n\Delta_{n}}\sum_{j=1}^{n}\mathbb{E}[(\zeta_{j}^{n})^{2}\mid\mathscr{G}_{j-1}^{n}]$ $\overset{\mathbb{P}}{\rightarrow}\sum_{k,k^{\prime}}z_{k}z_{k^{\prime}}\mathbb{E}(Z_{t_{k}}Z_{t_{k^{\prime}}}),$ (57)
@@
We verify relations (57) and (58) as follow:
Asymptotic Variance: In Lemmas 7 and 8 in Section A, it is shown that
@@
$\sigma_{n}^{2}=\Delta_{n}^{2}\sum_{l=0}^{n-1}\sum_{l^{\prime}=0}^{n-1}\overline{c}_{l}^{n}\overline{c}_{l^{\prime}}^{n}\left(\sum_{\ell,\ell^{\prime}=1}^{4}F_{n}^{(\ell,\ell^{\prime})}(t_{l},t_{l^{\prime}})\right)+\mathrm{o}_{\mathbb{P}}(1),$ (59)
where $F_{n}^{(\ell,\ell^{\prime})}$ is a sequence of bounded functions vanishing outside of $0,n\Delta_{n}]$ and such that (recall that $\tau_{n}(s)=[s/\Delta_{n}]\Delta_{n}$) $\sum_{\ell,\ell^{\prime}=1}^{4}F_{n}^{(\ell,\ell^{\prime})}(\tau_{n}(s),\tau_{n}(r))\rightarrow\Sigma_{a}(s,r)$, as $n\rightarrow\infty$. This means that if $c_{l}^{n}(t)=\mathbf{1}_{[t/\Delta_{n}]\leq l\leq n-1}$, then as $n\rightarrow\infty$, the continuity of $\dot{g}$ implies that
$\sigma_{n}^{2}\overset{\mathbb{P}}{\rightarrow}\sum_{\ell,\ell^{\prime}=1}^{4}\sum_{q,q^{\prime}=1}^{M}z_{q}z_{q^{\prime}}\int_{0}^{t_{q}}\int_{0}^{t_{q^{\prime}}}\dot{g}(a(s))\dot{g}(a(r))\Sigma_{a}(s,r)\mathrm{d}r\mathrm{d}s,$
which is exactly (57). On the other hand, if either $c_{l}^{n}(t)=\mathbf{1}_{[t/\Delta_{n}]\leq l\leq n-1}$ or $c_{l}^{n}(t)=\mathbf{1}_{[t/\Delta_{n}]\leq l\leq N_{n}-1}$
$\sigma_{n}^{2}=\sum_{\ell,\ell^{\prime}=1}^{4}\sum_{q,q^{\prime}=1}^{M}z_{q}z_{q^{\prime}}\int_{t_{q}}^{m_{n}\Delta_{n}}\int_{t_{q^{\prime}}}^{m_{n}\Delta_{n}}\dot{g}(a(\tau_{n}(s)))\dot{g}(a(\tau_{n}(r)))F_{n}^{(\ell,\ell^{\prime})}(t_{l},t_{l^{\prime}})\mathrm{d}r\mathrm{d}s,$ (60)
plus an error term that is asymptotically negligible. Therefore, in order to achieve (57) in this situation, it is enough to check that the (Generalized) Dominated Convergence Theorem (DCT from now on) can be applied in (60). To see that the latter is indeed possible, observe that from (11) we have that
$\rvert\dot{g}(a(\tau_{n}(s)))\rvert\lesssim\mathbf{1}_{s\leq\Delta_{n}}+a(s-\Delta_{n})\mathbf{1}_{s>\Delta_{n}}=:f_{n}(s)\rightarrow a(s),$ (61)
for all $s>0$. By the integrability and boundedness of $a$ we further have
$\int_{t}^{m_{n}\Delta_{n}}f_{n}(s)\mathrm{d}s=\int_{t}^{m_{n}\Delta_{n}}a(s)\mathrm{d}s+\mathrm{o}(1)\rightarrow\int_{t}^{\infty}a(s)\mathrm{d}s,\,\,\,t\geq 0.$ (62)
The previous two properties together with the boundedness of $F_{n}^{(\ell,\ell^{\prime})}$ justify the use of the Generalized DCT in (60).
@@
Lyapunov’s condition: Arguing as in the proof of Lemma 6 we get that $\Delta_{n}\sum_{l=0}^{j}\rvert\overline{c}_{l}^{n}\rvert$ is uniformly bounded. In consequence, if we choose $2<r<\frac{r_{0}}{2}\land 2\alpha$, we infer from Jensen’s inequality, Lemma 2, and Assumption 1 that for every $\ell=1,2,3,4$
$\mathbb{E}(\mid\zeta_{j}^{n,(\ell)}\mid^{r})\leq\left(4\Delta_{n}\sum_{l=0}^{j}\rvert\overline{c}_{l}^{n}\rvert\right)^{r-1}\Delta_{n}\sum_{l=0}^{j}\rvert\overline{c}_{l}^{n}\rvert\mathbb{E}(\mid\xi_{j,l,n}^{(\ell)}\mid^{r})\leq C\Delta_{n},$
from which (58) trivially holds.∎
Appendix A Computing the AVAR
In this subsection we show that the approximation (59) is valid. We will constantly use the estimates and identities presented in Lemma 1 as well as the basic fact that the variance of a sum of martingale differences is the sum of each individual variance. Thus, the reader must be aware that we will often omit referring to these results unless the context is not clear enough. Recall that $m_{n}=n$ or $m_{n}=N_{n}$, depending whether $c_{l}^{n}(t)=\mathbf{1}_{[t/\Delta_{n}]\leq l\leq n-1}$ or $c_{l}^{n}(t)=\mathbf{1}_{[t/\Delta_{n}]\leq l\leq N_{n}-1}$.

### v3

@@ 4.3. Proof of Theorems 5 and 6
which again implies $U^{n,1}\overset{u.c.p}{\to}0$.
Case $i=2$. Let $\tilde{m}_{n}=[T/\Delta_{n}]$ when $c_{l}^{n}(t)=\mathbf{1}_{0\leq l\leq[t/\Delta_{n}]-1}$ and $\tilde{m}_{n}=m_{n}$ otherwise. Using (42), we get that
$\sup_{t\in[0,T]}\rvert U_{t}^{n,2}\rvert\lesssim\sqrt{n\Delta_{n}^{3}}\int_{0}^{\tilde{m}_{n}\Delta_{n}}g^{\prime}(a(\tau_{n}(s)))\phi(s)\mathrm{d}s.$
Since $\sqrt{n\Delta_{n}^{3}}\rightarrow 0,$ and $\phi$ is bounded (being càdlàg and vanishing at $\infty$), it suffices to show that $\int_{0}^{m_{n}\Delta_{n}}g^{\prime}[a(\tau_{n}(s))]\mathrm{d}s$ is bounded over $n$. If $\tilde{m}_{n}=[T/\Delta_{n}]$ this is immediate from the continuity of $g^{\prime}$. If $\tilde{m}_{n}=m_{n}$ and $g\in\mathfrak{C}_{p,q}^{2}$, Remark 1 implies that
$\int_{0}^{m_{n}\Delta_{n}}\rvert g^{\prime}[a(\tau_{n}(s))]\rvert\mathrm{d}s\lesssim\int_{0}^{m_{n}\Delta_{n}}a(\tau_{n}(s))^{p+1}\mathrm{d}s\lesssim\Delta_{n}+\int_{0}^{\infty}a(s)^{p+1}\mathrm{d}s,$ (55)
just as required.
Case $i=3$. Since $xa(x)\rightarrow 0$, as $x\rightarrow\infty$,
$\sup_{t\in[0,T]}\rvert U_{t}^{n,3}\rvert\lesssim_{a}\frac{1}{\sqrt{n\Delta_{n}}}\int_{0}^{\tilde{m}_{n}\Delta_{n}}g^{\prime}[a(\tau_{n}(s))]\mathrm{d}s\rightarrow 0,$
due to (55).
Case $i=4$. Using the same arguments as above for $i=4$ along with the the stationarity of $X$, we obtain $\sup_{t\in[0,T]}|U_{t}^{n,4}|\overset{\mathbb{P}}{\to}0.$
Case $i=5$. We first consider the case $c_{l}^{n}(t)=\mathbf{1}_{0\leq l\leq[t/\Delta_{n}]-1}$. Since $g\in\mathcal{C}^{2}$ and $a$ is Lipschitz (thanks to Assumption 1), we deduce as in (55)
$\rvert U_{t}^{n,5}\rvert\lesssim\sqrt{n\Delta_{n}}\int_{0}^{[t/\Delta_{n}]\Delta_{n}}\left|g(a(\tau_{n}(s)))-g(a(s))\right|\mathrm{d}s+\sqrt{n\Delta_{n}^{3}}\lesssim\sqrt{n\Delta_{n}^{3}}\rightarrow 0.$
Assume now $c_{l}^{n}(t)=\mathbf{1}_{[t/\Delta_{n}]\leq l\leq m_{n}-1}$ and let $g\in\mathfrak{C}_{p,q}^{2}$. Then,
$\rvert g(a(\tau_{n}(s)))-g(a(s))\rvert\lesssim\Delta_{n}a(\tau_{n}(s)),$
by the Mean-Value Theorem, the Lipschitz property of $a$, and the fact that $g^{\prime}(x)=\mathrm{o}(|x|)$ as $x\to 0$. Furthermore, from (11) we also have that $\rvert g(a(s))\rvert\leq Ca(s)^{2+p}$. Thus,
$\rvert\mathcal{R}_{t}^{n}\rvert\lesssim\Delta_{n}\int_{0}^{m_{n}\Delta_{n}}a(\tau_{n}(s))\mathrm{d}s+\int_{m_{n}\Delta_{n}}^{\infty}a(s)^{2+p}\mathrm{d}s+\Delta_{n}.$
Using this, Assumption 1, and (55) we obtain that, uniformly on $t\in[0,T]$,
@@
Therefore, if $m_{n}=n$, we have $\sqrt{n\Delta_{n}}(m_{n}\Delta_{n})^{1-\alpha(p+2)}=(n\Delta_{n})^{\frac{3}{2}-\alpha(p+2)}\rightarrow 0$ because $\alpha>1$. If instead $m_{n}=N_{n}$, Assumption 2 and (16), again imply
@@
The next step is to write $V^{n}$, defined in (53), as a sum of martingale differences plus an error term. Once this is achieved, the proof of Theorems 5 and 6 will follow as an application of, for instance, Theorem 6.1 in [9] (c.f. Theorem IX 7.19 in [13]). To do this, we use the decomposition (32) and write
@@
where (see (47))
$V_{t}^{\prime n}$= $\sqrt{\frac{\Delta_{n}}{n}}\sum_{l=0}^{n-1}\sum_{j=l}^{n-1}c_{l}^{n}(t)g^{\prime}(a(l\Delta_{n}))\varXi_{j,l,n};$ (56)
$U_{t}^{\prime,n}$= $\sqrt{\frac{\Delta_{n}}{n}}\sum_{l=0}^{n-1}\sum_{j=l}^{n-1}c_{l}^{n}(t)g^{\prime}(a(l\Delta_{n}))R_{j,l,n}.$
Note that $\varXi_{j,l,n}$ is $\mathscr{G}_{j}^{n}$-measurable (see (23)) and, by construction, $\mathbb{E}(\varXi_{j,l,n}\rvert\mathscr{G}_{j-1}^{n})=0$, so $(\varXi_{j,l,n}:0\leq l\leq j)_{j\geq 0}$ forms a martingale difference sequence with respect to $(\mathscr{G}_{j}^{n})_{j\geq 0}$. The next lemma shows that $U^{\prime,n}$ is asymptotically negligible.
@@
As before, we may assume without loss of generality that $L$ has mean zero. Since each $\rho_{j,l,n}^{(\ell)}$ (see (34)) is a martingale difference, we only need to show that for $\ell=1,2,3$
$\frac{1}{n\Delta_{n}}\mathcal{E}_{n}^{(\ell)}(t):=\frac{1}{n\Delta_{n}}\sum_{j=0}^{n}\mathbb{E}\left[\left(\Delta_{n}\sum_{l=0}^{j}c_{l}^{n}(t)g^{\prime}(a(l\Delta_{n}))\rho_{j,l,n}^{(\ell)}\right)^{2}\right]\rightarrow 0.$
To see this, first note that the quantity $\Delta_{n}\sum_{l=0}^{j}\rvert c_{l}^{n}(t)g^{\prime}(a(l\Delta_{n}))\rvert$ is uniformly bounded. Indeed, if $c_{l}^{n}(t)=\mathbf{1}_{0\leq l\leq[t/\Delta_{n}]-1}$, the continuity of $g^{\prime}$ and $a$ implies
$\Delta_{n}\sum_{l=0}^{j}\rvert c_{l}^{n}(t)g^{\prime}(a(l\Delta_{n}))\rvert\leq\Delta_{n}\sum_{l=0}^{[T/\Delta_{n}]}\lvert g^{\prime}(a(l\Delta_{n}))\rvert\leq CT.$
If instead $c_{l}^{n}(t)=\mathbf{1}_{[t/\Delta_{n}]\leq l\leq n-1}$ or $c_{l}^{n}(t)=\mathbf{1}_{[t/\Delta_{n}]\leq l\leq N_{n}-1}$ and $g\in\mathfrak{C}_{p,q}^{2}$, the same bound follows from (55). Hence, by Jensen’s inequality,
$\mathcal{E}_{n}^{(\ell)}(t)\lesssim\Delta_{n}\sum_{j=0}^{n}\sum_{l=0}^{j}\rvert c_{l}^{n}(t)g^{\prime}(a(l\Delta_{n}))\rvert\mathbb{E}[(\rho_{j,l,n}^{(\ell)})^{2}].$
Moreover, Lemma 1 yields
@@
and, by the Dominated Convergence Theorem,
@@
We are now ready to prove the CLTs for our statistics. In view of the previous two lemmas and the definition of $\tilde{a}$ (see (26)), we may and do assume that $L$ has mean 0. Using the decomposition $V_{t}^{n}=V_{t}^{\prime n}+U_{t}^{\prime,n}$, the proof reduces to applying a martingale central limit theorem to the leading term $V^{\prime n}$.
@@
We remind the reader that $\mathcal{S}^{n}(g)$ and $\mathcal{S}(g)$ denote any of the pairs $(\Psi^{n}(g),\Psi(g))$, $(\Lambda^{n}(g),\Lambda(g))$, or $(\bar{\Lambda}^{n}(g),\bar{\Lambda}(g))$. Recall also the decomposition
$\sqrt{n\Delta_{n}}\left(\mathcal{S}_{t}^{n}(g)-\mathcal{S}_{t}(g)\right)=V_{t}^{n}+\sum_{\ell=1}^{5}U_{t}^{n,\ell}.$
By Lemma 5 and Remark 2, it is enough to show that $V^{n}\rightarrow Z$ in $\mathcal{D}([0,T])$, where $Z$ is the limit process stated in the theorems. We proceed in three steps. First we establish tightness of $V^{n}$. Next we identify the finite-dimensional limits using the Cramér-Wold device. Finally we verify the martingale CLT conditions.
Step 1: Tightness. From Lemma 3, we deduce that for all $0\leq s\leq t\leq T$
@@
Therefore, by Theorem 13.5 and equation (13.14) in [5], the sequence $V^{n}$ is tight on $\mathcal{D}([0,T])$.
Step 2: Cramér-Wold reduction. Fix $M\in\mathbb{N}$, $0=t_{0}\leq t_{1}\leq\cdots\leq t_{M}\leq T$ , and $z_{1},z_{2},\ldots,z_{M}\in\mathbb{R}$. By Lemma 6, convergence of the finite-dimensional distributions of $V^{n}$ towards those of $Z$ reduces to showing $\sum_{k=1}^{M}z_{k}V_{t_{k}}^{\prime n}\overset{d}{\rightarrow}\sum_{k=1}^{M}z_{k}Z_{t_{k}}$, with $V^{\prime n}$ as in (56). Define
$\overline{c}_{l}^{n}=\sum_{k=1}^{M}z_{k}c_{l}^{n}(t_{k})g^{\prime}(a(l\Delta_{n})),\quad\zeta_{j}^{n}=\Delta_{n}\sum_{l=0}^{j}\overline{c}_{l}^{n}\varXi_{j,l,n},$ (57)
in which $\varXi_{j,l,n}=\sum_{\ell=1}^{4}\xi_{j,l,n}^{(\ell)}$ (see (33)). Then
@@
Since $\zeta_{j}^{n}$ is a $(\mathscr{G}_{j}^{n})$-martingale difference (see our discussion before Lemma 6), the CLT of Theorem 6.1 in [9] (c.f. Theorem IX 7.19 in [13]) applies provided that for some $r>2$
$1.\,\sigma^{2}_{n}:=\frac{1}{n\Delta_{n}}\sum_{j=1}^{n}\mathbb{E}[(\zeta_{j}^{n})^{2}\mid\mathscr{G}_{j-1}^{n}]$ $\overset{\mathbb{P}}{\rightarrow}\sum_{k,k^{\prime}}z_{k}z_{k^{\prime}}\mathbb{E}(Z_{t_{k}}Z_{t_{k^{\prime}}}),$ (58)
@@
Step 3: Verification of (58) and (59).
Asymptotic Variance: In Lemmas 7 and 8 in Appendix A, it is shown that
@@
$\sigma^{2}_{n}=\Delta_{n}^{2}\sum_{l=0}^{n-1}\sum_{l^{\prime}=0}^{n-1}\overline{c}_{l}^{n}\overline{c}_{l^{\prime}}^{n}\left(\sum_{\ell,\ell^{\prime}=1}^{4}F_{n}^{(\ell,\ell^{\prime})}(t_{l},t_{l^{\prime}})\right)+\mathrm{o}_{\mathbb{P}}(1),$ (60)
where $F_{n}^{(\ell,\ell^{\prime})}$ are bounded functions vanishing outside $0,n\Delta_{n}]$ and satisfying
$\sum_{\ell,\ell^{\prime}=1}^{4}F_{n}^{(\ell,\ell^{\prime})}(\tau_{n}(s),\tau_{n}(r))\rightarrow\Sigma_{a}(s,r),\text{ as }n\rightarrow\infty.$
Thus, if $c_{l}^{n}(t)=\mathbf{1}_{[t/\Delta_{n}]\leq l\leq n-1}$, the continuity of $g^{\prime}$ immediately yields
$\sigma^{2}_{n}\rightarrow\sum_{\ell,\ell^{\prime}=1}^{4}\sum_{q,q^{\prime}=1}^{M}z_{q}z_{q^{\prime}}\int_{0}^{t_{q}}\int_{0}^{t_{q^{\prime}}}g^{\prime}(a(s))g^{\prime}(a(r))\Sigma_{a}(s,r)\mathrm{d}r\mathrm{d}s,$
which coincides with (58). If instead $c_{l}^{n}(t)=\mathbf{1}_{[t/\Delta_{n}]\leq l\leq n-1}$ or $c_{l}^{n}(t)=\mathbf{1}_{[t/\Delta_{n}]\leq l\leq N_{n}-1}$
$\sigma^{2}_{n}=\sum_{\ell,\ell^{\prime}=1}^{4}\sum_{q,q^{\prime}=1}^{M}z_{q}z_{q^{\prime}}\int_{t_{q}}^{m_{n}\Delta_{n}}\int_{t_{q^{\prime}}}^{m_{n}\Delta_{n}}g^{\prime}(a(\tau_{n}(s)))g^{\prime}(a(\tau_{n}(r)))F_{n}^{(\ell,\ell^{\prime})}(t_{l},t_{l^{\prime}})\mathrm{d}r\mathrm{d}s+\mathrm{o}_{\mathbb{P}}(1),$ (61)
Hence, to achieve (58), it suffices to justify the use of the generalized dominated convergence theorem. To see this is the case, observe that from (11) we have that
$\rvert g^{\prime}(a(\tau_{n}(s)))\rvert\lesssim\mathbf{1}_{s\leq\Delta_{n}}+a(s-\Delta_{n})\mathbf{1}_{s>\Delta_{n}},$ (62)
and the right-hand side of this converges pointwise to $a$ and is bounded. Hence the generalized dominated convergence applies.
@@
Lyapunov’s condition: Arguing as in the proof of Lemma 6, the quantity $\Delta_{n}\sum_{l=0}^{j}\rvert\overline{c}_{l}^{n}\rvert$ is uniformly bounded. Choose $2<r<\frac{r_{0}}{2}\land 2\alpha$. Then Jensen’s inequality, Lemma 2, and Assumption 1 imply that
$\mathbb{E}(\mid\zeta_{j}^{n,(\ell)}\mid^{r})\lesssim\Delta_{n}\sum_{l=0}^{j}\rvert\overline{c}_{l}^{n}\rvert\mathbb{E}(\mid\xi_{j,l,n}^{(\ell)}\mid^{r})\lesssim C\Delta_{n},$
Hence, condition (59) holds concluding this the proof. ∎
[…]
Appendix A Approximations of the Asymptotic Variance
In this subsection we justify the approximation (60). Throughout the analysis we repeatedly use the moment estimates and identities from Lemma 1, as well as the fact that the variance of a sum of martingale differences equals the sum of the individual variances. To avoid unnecessary repetition, these arguments will not be explicitly referenced each time they are applied. Recall that $m_{n}=n$ or $m_{n}=N_{n}$, depending on whether $c_{l}^{n}(t)=\mathbf{1}_{[t/\Delta_{n}]\leq l\leq n-1}$ or $c_{l}^{n}(t)=\mathbf{1}_{[t/\Delta_{n}]\leq l\leq N_{n}-1}$.